OpenStudy (anonymous):

solve the system of equations. x-4y=5 x=3y-2 (__,__)

OpenStudy (muzzack):

wait is x-4 and y=5 seperate

OpenStudy (anonymous):

x-4y=5<-eqn1 x=3y-2<- eqn 2 rearrange the eqn 2

OpenStudy (anonymous):

x - 3y = -2

OpenStudy (adamk):

Once you rearrange equation 2, subtract the all the terms.

OpenStudy (muzzack):

oh got it

OpenStudy (adamk):

So you're left with -y = 7 or y = -7

OpenStudy (muzzack):

adamk explain how to use the elimination

OpenStudy (anonymous):

so its (7, -7) ?

OpenStudy (adamk):

Then, substitute -7 in for "y" in one of the equations to get "x". Then, you have your ordered pair, (x, y).

OpenStudy (adamk):

Wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sorry thats another question i need help on

OpenStudy (adamk):

The ordered pair is (-23, -7). Go look up systems of equations on khan academy. It's difficult to explain here.

OpenStudy (anonymous):

k thanks